Alan, Betty, and Cathy each have the same amount of money. How much money should Betty give to Alan and Cathy so that Alan has $5 more than Betty, and Cathy has $1 more than Alan?
Answer by ikleyn(52908)

.
Let M be the amount of money each person had initially.

Let A = amount Betty gave to Alan

Let C = amount Betty gave to Cathy.


Then Betty has  (M-A-C).
     Alan  has  (M+A).
     Cathy has  (M+C).


And the condition says


M - A - C = M + A - 5     (1)    ( . . . so that Alan has $5 more than Betty)
M + C     = M + A - 1     (2)    ( . . . Cathy has $1 more than Alan)


Simplify by canceling M:


-A -C = A - 5             (1')
C     = A - 1             (2')


Write in the standard form

2A + C = 5                (1'')
 A - C = 1                (2'')


Now add eq(1'')  and eq(2'').  You will get

3A = 6  ====>  A = 2.


Then from equation (2'')  C = A - 1 = 1.


Answer.   A = 2, C = 1.  In other words,  B gives 2 to A and B gives 1 to C.


Check.    Do the check on your own.


------------------
Comment from student:   Thank you for  #1112120
------------------


My response:   I want to commend you for two things:

            a)  for sending your "Thanks",   and

            b)  for referring to the problem ID number.


Most of other users/"students" are so badly educated people that allow themselves do neither  a)  nor  b).